# Attention U-Net for all-sky continuous gravitational wave searches

### v1

@@ IV.4 Sky position
Here, we investigate the sensitivity as a function of the CW source sky position. To obtain the intrinsic dependence on sky position, we fixed the SNR of the injected signal by adjusting the noise level to remove the dependence on the detector antenna patterns (Dreissigacker and Prix, 2020):
@@
where $R^{2}(\theta)$ is a geometric factor that depends on $\{\alpha,\delta,\cos\iota,\psi\}$. We fixed the SNR so that the overall $p_{\mathrm{det}}\approx 50\%$.
Figure 7: Detection probability $p_{\mathrm{det}}$ (at a fixed $p_{\mathrm{fa}}=1\%$ per $50\,\mathrm{mHz}$) as a function of the CW source sky position (Hammer projection). The SNR is fixed to have an overall $p_{\mathrm{det}}\approx 50\%$. The top panel shows the all-sky detectability at 20 Hz (left) and 500 Hz (right). The bottom panel shows the all-sky detectability at 1000 Hz (left) and 20–1000 Hz (right).
In Figure. 7, we observe a clear preference in the sensitivity for the 20 and 500,Hz models (top panel). They are less sensitive near the celestial equator and more sensitive for the CW signals coming from the equatorial poles. The 1000 Hz model (bottom left) shows no clear preference toward a certain direction, but an irregular pattern of the sensitivity. The model trained on the entire 20–1000 Hz band (bottom right) shows a mixture of these sensitivity patterns. (Dreissigacker and Prix, 2020) and (Joshi and Prix, 2023) pointed out that the CNN-based neural network struggles to learn the structure of a comparatively wider signal due to their larger Doppler modulation. Our results align with this, indicating that the Attention U-Net is less sensitive to signals coming from near the celestial equator ($\delta\approx 0$), where Doppler modulation is larger, and at higher frequencies, where the signal’s frequency span increases proportionally with frequency.
@@ V Discussion
The attention U-Net used in this work effectively denoises and detects CW signals in noisy data, leveraging a distinct architecture compared to existing methods. By employing a normalized loss function (Eq. 5) that balances signal restoration and noise suppression, along with a detection statistic based on average power across image dimensions, the model successfully recovers CW signals, as shown in Figures 2 and 3. The attention U-Net shows comparable sensitivity to that reported in (Joshi and Prix, 2024). While (Joshi and Prix, 2024) offers a better high-frequency sensitivity, our architecture balances performance and efficiency as a promising alternative. The model demonstrates robustness to data time gaps, with sensitivity $\propto\eta^{0.53\pm 0.02}$ with duty factor. The model’s sensitivity scales with the total timespans of the data as $\propto T^{0.28\pm 0.01}$, while the training time scales as $\sim T^{m}$, where $1<m<2$. Sky position analysis reveals frequency-dependent sensitivity variations, with higher detectability near equatorial poles at lower frequencies and more irregular sensitivity patterns at higher frequencies, indicating challenges due to broader Doppler modulations.
@@
To enhance performance, we attempted curriculum learning by starting with less noisy data and gradually increasing the noise level, but it yielded no clear sensitivity gains. However, scaling up to 10,000 training samples and 128 initial feature channels improved $\mathcal{D}^{90\%}$ to $18.00\pm 0.12\,\mathrm{Hz}^{-1/2}$ at 1000 Hz, suggesting optimization potential. Due to limited resources, we show the results as a proof-of-concept, showing U-Net-based architectures are viable for all-sky CW searches over wide parameter spaces, rather than to train the model with more data or channels. For addressing Doppler modulation challenges, increasing convolutional kernel sizes (e.g., to $\times4$ or $\times5$) was tested but led to unstable training and reduced sensitivity, warranting further investigation into effective capture of larger signal patterns. Similarly, integrating residual blocks—as successfully applied in (Dreissigacker and Prix, 2020) and (Joshi and Prix, 2024)—through a Residual U-Net (Alom et al., 2018) was attempted, but resulted in unstable training and lower sensitivity compared to the Attention U-Net. Additional studies are needed to study the reason behind it.
@@
Looking ahead, future work could include fine-tuning high-frequency models with larger datasets, incorporate time gaps in training for greater robustness, or replace MaxPooling with wavelet transforms (Li et al., 2020) to better preserve weak signals. Evaluating non-stationary noise in long observations is also essential for real data applications. The neural network could be applied to all-sky searches for binary NSs by training on data preprocessed using the TwoSpect method (Goetz and Riles, 2011), which transforms long-duration signals into localized patterns that may be less affected by the kernel size limitations encountered here. The attention U-Net’s output could also serve as input to a variational autoencoder (Gabbard et al., 2021) for rapid parameter estimation. Finally, having demonstrated the capability of the U-Net architecture to detect CW signals, exploring diffusion models (Ho et al., 2020; Song et al., 2021; Legin et al., 2025)—advanced denoising frameworks that employ U-Net as a backbone—may further enhance the sensitivity of machine learning–based CW searches.

### v2

@@ IV.4 Sky position
In this section, we investigate the sensitivity of the all-sky search as a function of the CW source’s sky position. To obtain the intrinsic dependence on sky position, we fixed the SNR of the injected signal by adjusting the noise level, thereby removing the influence of detector antenna patterns (Dreissigacker and Prix, 2020), as given by:
@@
where $R^{2}(\theta)$ is a geometric factor that depends on $\{\alpha,\delta,\cos\iota,\psi\}$. By fixing the SNR to achieve an overall $p_{\mathrm{det}}\approx 50\%$ across all sky positions, we calculate the overall $p_{\mathrm{det}}$ for various sky positions using the noise level derived from Eq. 7.
Figure 7: Detection probability $p_{\mathrm{det}}$ (at a $1\%$ false alarm rate per $50\,\mathrm{mHz}$) of the attention U-Net models as a function of the CW source sky position (Hammer projection). The SNR is fixed to have an overall $p_{\mathrm{det}}\approx 50\%$. The top panel shows the all-sky detectability for models trained at 20 Hz (left) and 500 Hz (right). The bottom panel shows the all-sky detectability for models trained at 1000 Hz (left) and 20–1000 Hz (right).
[…]
@@ V Discussion
The attention U-Net used in this work effectively denoises and detects CW signals in noisy data, leveraging a distinct architecture compared to existing methods. By employing a normalized loss function (Eq. 5) that balances signal restoration and noise suppression, along with a detection statistic based on mean power across image dimensions, the model successfully recovers CW signals, as shown in Figures 2 and 3. The attention U-Net shows comparable sensitivity to that reported in (Joshi and Prix, 2024). While (Joshi and Prix, 2024) offers a better high-frequency sensitivity, our architecture balances performance and efficiency as a promising alternative. The model demonstrates robustness to data time gaps, with sensitivity $\propto\eta^{0.53\pm 0.02}$ with duty factor. The model’s sensitivity scales with the total timespans of the data as $\propto T^{0.28\pm 0.01}$, while the training time scales as $\sim T^{m}$, where $1<m<2$. Sky position analysis from Figure 7 reveals frequency-dependent sensitivity variations, with higher detectability away from the celestial equator at lower frequencies and more irregular sensitivity patterns at higher frequencies, indicating challenges due to broader Doppler modulations.
@@
To enhance performance, we attempted curriculum learning by starting with less noisy data and gradually increasing the noise level, but it yielded no clear sensitivity gains. However, scaling up to 20,000 training samples (10,000 signals + 10,000 pure noise) and 128 initial feature channels improved $\mathcal{D}^{90\%}$ to $18.00\pm 0.12\,\mathrm{Hz}^{-1/2}$ at 1000 Hz, suggesting optimization potential. Due to limited resources, we show the results as a proof-of-concept, showing U-Net-based architectures are viable for all-sky CW searches over wide parameter spaces, rather than to train the model with more data or channels. For addressing Doppler modulation challenges, increasing convolutional kernel sizes (e.g., to $\times4$ or $\times5$) was tested but led to unstable training and reduced sensitivity, warranting further investigation into effective capture of larger signal patterns. Similarly, integrating residual blocks—as successfully applied in (Dreissigacker and Prix, 2020) and (Joshi and Prix, 2024)—through a Residual U-Net (Alom et al., 2018) was attempted, but resulted in unstable training and lower sensitivity compared to the attention U-Net. Additional studies are needed to study the reason behind it.
@@
Looking ahead, future work could include fine-tuning high-frequency models with larger datasets, incorporate time gaps in training for greater robustness, or replace MaxPooling with wavelet transforms (Li et al., 2020) to better preserve weak signals. Evaluating non-stationary noise in long observations is also essential for real data applications. The neural network could be applied to all-sky searches for CWs emitted from binary NSs by training on data preprocessed using the TwoSpect method (Goetz and Riles, 2011), which transforms long-duration signal spectrograms into localized patterns that may be less affected by the kernel size limitations encountered here. The attention U-Net’s output could also serve as input to a variational autoencoder (Gabbard et al., 2021) for rapid parameter estimation. Finally, having demonstrated the capability of the U-Net architecture to detect CW signals, exploring diffusion models (Ho et al., 2020; Song et al., 2021; Legin et al., 2025)—advanced denoising frameworks that employ U-Net as a backbone—may further enhance the sensitivity of machine learning–based CW searches.